Sir, i got clear with VIT but didn’t got much good rank......should i proceed for counselling or not...?????

Sir,
i got clear with VIT but didn’t got much good rank......should i proceed for counselling or not...?????

Grade:12th pass

36 Answers

Shubham Singh
8 years ago
if u can tell the rank then i can suggest anything
mansi dabriwal
8 years ago
it is 87005...
Shubham Singh
8 years ago
u should not. the closing ranks are within 20000.
my rank is 5045 and i m not sure whether i will get desired branch.
mansi dabriwal
8 years ago
r u going for the counselling?????????
VAISHNAVI BATHINA
8 years ago
you cannot get into vit with 87005 rank so it is best not to proceed for counselling
 
Shubham Singh
8 years ago
yes i m going for counselling
mansi dabriwal
8 years ago
i think ur counseling might have finished.....so have u got ur desired branch??????
Shubham Singh
8 years ago
yes, i got my desired branch and i got information there that if someone leaves college after getting a seat, then his seat will be left vacant
mansi dabriwal
8 years ago
that means no new student will be admitted to that place....???
ng29
8 years ago
hey mansi
just leave the dreams for VIT , and prepare for bitsat 
and have u done cd of arihant bitsat , whats ur average score
 
approve if useful
mansi dabriwal
8 years ago
i m doing the papers gvn in book itself...actually my main mistake is that i have opted for any couching during my 11th and12th class....
ng29
8 years ago
i does not understand ur previous statement
that whether u have opted or not for coaching
it matters a lot as some of students start preparing since 9 or 10 class
but dont worry prepare for bitsat and upes 
all the best
 
approve if useful
mansi dabriwal
8 years ago
up’s sorry ...i din’t make that clear...actually i hv never opted for couching n now i regret for it... 
ng29
8 years ago
hmmmm. now i understand
but dont worry and stay calm 
let me ask u question logical reasoning
if TREE IS CODED AS 91 
FEED IS CODED AS 47 THEN WHAT IS CODE FOR MEET
answer it and reply
 
approve if useful
mansi dabriwal
8 years ago
86...???? if my ans. is wrong  thn could u explain your logic....????
mansi dabriwal
8 years ago
plz reply fast eager to know the answer....
mansi dabriwal
8 years ago
can i get the ans. plzzzzzz....
ng29
8 years ago
no answer is 118
ng29
8 years ago
 
there is a pattern in this question
TREE               20*1 + 18*2 + 5*3 5*4 = 20+36+15+20 = 91  
FEED               6*1 + 5*2 + 5*3 + 4*4 = 6+10+15+16 = 47
MEET             13*1 +  5*2  +  5*3  +  20*4  = 13+10+15+80 =118
i will explain u for TREE 
RANKING OF T IN ALPHABETICAL ORDER IS 20     AND  1  PLACE IN WORD
RANKING OF R IN ALPHABETICAL ORDER IS 18       AND  2  PLACE IN WORD
RANKING OF E IN ALPHABETICAL ORDER IS 5           AND  3  PLACE IN WORD
RANKING OF E IN ALPHABETICAL ORDER IS 5              AND  4  PLACE IN WORD
 
and code is given as sum of multiplication of alphabet places in alphabetical order and in given word order
 
APPROVE IF USEFUL 
APPROVE IF USEFUL 
APPROVE IF USEFUL 
mansi dabriwal
8 years ago
gud ques.
ng29
8 years ago
i know it thats why i asked so that u can tackle these type of questions in bitsat
mansi dabriwal
8 years ago
k...if u hv more such ques. u can forward me....n it will really help me....
ng29
8 years ago
definitely i will post on our any of discussion
 
ng29
8 years ago
ok then try this question and dont forget to approve if u like the answer
 
 puzzle
 
A lady buys goods worth rs.200 from a shop. (shopkeeper selling the goods with zero profit).
The lady gives him 1000 rs note. The shopkeeper gets the change from the next shop and keeps 200 for himself and returns rs.800 to d lady.
Later the shopkeeper of the next shop comes with the 1000rs note saying "duplicate" and takes his money back.
"How much LOSS did the shopkeeper face ?"
mansi dabriwal
8 years ago
Rs.1000 is total loss of the shopkeeper.....
ng29
8 years ago
congrtats , it is correct answer 
u understood the logic na ??
mansi dabriwal
8 years ago
yah..thanx...according to my logic....the real loss of shopkeeper was when he gave back Rs.1000 in place of duplicate note is that so???? 
ng29
8 years ago
no  mansi, 
 
as lady get product of Rs200 + Rs800 cash so she is in profit of Rs1000
on the other hand poor shopkeeper suggered A LOSS of Rs1000
 
approve if useful
Ayushmaan Vardhan
8 years ago
You can check the cut off marks list on internet by visiting the website of VIT, if it is favourable then you must proceed otherwise you can think of some other reputed institutuions – for eg. Galgotia uni, LPU, Phagwara, BHU, BITS pilani
mansi dabriwal
8 years ago
but according to ur logic ng29 the shopkeeper will suffer a loss of 2000 as 1000 due to lady and 1000 when he returned back 1000 to the neighbouring shop whomm he hv exchanged the note as he was not having change of 1000
ng29
8 years ago
no mansi u r comiitting mistake and dos not inderstand the situation comletely
try again to understand it , and focus on dis staement
as lady get product of Rs200 + Rs800 cash so she is in profit of Rs1000
 
ng29
8 years ago
do u get info about ur result
are u giving jee advance on 24 may
 
approve if useful (all)
mansi dabriwal
8 years ago
no info till now...many are saying it will b declared in july...don noe what will happen....no i m not eligible for advance as i hv not cleared paper-1 of mains...
ng29
8 years ago
dont worry and relax
its destiny and all will be ok, 
mansi dabriwal
8 years ago
yah...now i hv kept everything on my destiny only....
Ayushmaan Vardhan
8 years ago
If not at VIT then try at other places like JJTU, Rajasthan, DTU Delhi, JJTU Jhunjhunu, Lovely Professional University, Phagwara